INAF queue observing, 2015 Nov 28-29

Observer: F. Cusano, A. Rossi
Telescope Operator: Steve Allanson
Support Astronomer: Dave Thompson (on call)
Instrument: LUCI/MODS


At the sunset the sky is clear. The wind is borderline (16 m/s).

Night duration (12\xB0 el twilight) 18:20-6:00 ( 02:30-11:50 UT) = 11 h 40m

Observing time: 09 hrs 50 m (84%)

Weather time loss: 0 m (0%)*

technical time loss 1h 50m (16%)

Data Summary

Project_RUN Objects Filter ExpTime Notes
Ragazzoni PY_340 210zJHK 30m (weak alignement for LUCI mos problem)
Saracco M004XLSms2 Red grating 1h45m  
Standard GD71 red grating 9m  
Guaita M0717 H2 1h another 30 min have been done the 20151125 during LBTO service
Guaita M0717 Br_Gam 1h  
Guaita M0717 Ks 10m  
Standard Rubin149 H2 10m  
Standard P266-C H2, Br_gam, Ks 10m,10m,2m  
sky flat LUCI J,Ks, K, H2, Br_Gam    


00:00 problems with executing the OB for skyflat with LUCI
we cannot move the blind mask out because of connection and configuration problem.

00:34 skyflat in K band with LUCI2 luci2.20151129.0002-0014
  • skyflat J luci2.20151129.0015-0026

01:20 slewing to Ragazzoni PY_340_mod
  • seeing 0.7
  • airmass 1.1
  • acquisition, problem in the acquisition in moving the mask, problem solved changing the mask id in the ob, luci2.20151129.0027-0030
  • 1:50 science==luci2.20151129.0031-0035==
  • 1:50 seeing 1.3"

02:32 slewing to the telluric Ragazzoni
  • acquisition luci2.20151129.0036-0052
    problem executing OB, in the first attempt the star is outside the slit. A second attempt was succesfull

3:19 slewing to the telluric Iovino
  • acquisition luci2.20151129.0053-65

3:46 slewing to Iovino 2015_2016_5_HK_spec_251115
  • acquisition luci2.20151129.0066-0069
  • problems to align mask

4:30 swich to MODS for problem with LUCI2

4:47 slew to Saracco, acquisition mods4.XLSm2.acq mods2r.20151129.0002-0006
  • seeing 1"

5:08 execMods2 mods.4.XLSms2_1.obs mods2r.20151129.0007-10
  • at the start of exposure seeing 1.1"
  • at 3rd exposure seeing 0.9" (5:58)
  • at 4th exposure seeing 0.8" (6:15)

6:18 execMods2 mods.4.XLSms2_2.obs mods2r.20151129.0011-13
  • at the start of exposure seeing 0.9"
  • at 3rd exposure seeing 0.8" (6:51)

7:18 slew to standard GD71 for Saracco, acqMods mods2r.20151129.0014-0015
  • execMODS mods2r.20151129.0016-0018

7:50 swith to LUCI

8:03 slew to Guaita photometric standard Rubin149_H2, luci2.20151129.0070-0079
  • seeing 1.0"

08:12 slew to Guaita imaging H2, luci2.20151129.0080-139
  • seeing 0.8"

10:05 Guaita imaging Br_Gam luci2.20151129.140-199, problem with the guide star, the TO check the pointing (15min lost)
  • file luci2.20151129.140.fits not good
  • seeing 1"
  • at 10:36 seeing 0.7"

11:49 Guaita Ks luci2.20151129.200-209
  • seeing 0.8"

12:06 slewing to a standard, P266-C; no guide star, the TO is checking the pointing
  • 12:24 H2 luci2.20151129.210-219
  • Br_Gam luci2.20151129.221-230
  • Ks luci2.20151129.231-241

13:14 starting sky flat Ks, J, H2, Br_Gam

14:10 starting LUCI calibration Ragazzoni, PY_340
  • execMods luci2.20151129.328-336
Topic revision: r7 - 20 Dec 2017, RobertaCarini
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback